### v1

# Temporal Triplane Transformers as Occupancy World Models

Abstract
Recent years have seen significant advances in world models, which primarily focus on learning fine-grained correlations between an agent’s motion trajectory and the resulting changes in its surrounding environment. However, existing methods often struggle to capture such fine-grained correlations and achieve real-time predictions. To address this, we propose a new 4D occupancy world model for autonomous driving, termed T3Former. T3Former begins by pre-training a compact triplane representation that efficiently compresses the 3D semantically occupied environment. Next, T3Former extracts multi-scale temporal motion features from the historical triplane and employs an autoregressive approach to iteratively predict the next triplane changes. Finally, T3Former combines the triplane changes with the previous ones to decode them into future occupancy results and ego-motion trajectories. Experimental results demonstrate the superiority of T3Former, achieving 1.44$\times$ faster inference speed (26 FPS), while improving the mean IoU to 36.09 and reducing the mean absolute planning error to 1.0 meters.
Keywords Occupancy world models  $\cdot$ Triplane  $\cdot$ Multi-scale  $\cdot$ Autoregression
1 Introduction
footnotetext: Corresponding authors
World models [1, 2] are designed to predict future scenes and facilitate motion planning for agents. These models first construct lower-dimensional representations of the scenes, which serve as a foundation for learning the patterns of environmental dynamics. This capability supports the identification of potential dangers, the determination of traffic participants’ intentions, and ultimately leads to improved decision-making.
This paper focuses on world models for autonomous driving [3, 4, 5, 6, 7], where accurately predicting the future behavior of traffic participants is essential for the agent’s planning. Existing methods [8, 6, 7, 9] mainly provide instance-level predictions for traffic participants from a Bird’s Eye View (BEV) perspective, or directly utilize diffusion models [10, 11, 12, 13, 14] to generate future pixel-level driving views. However, these methods have difficulty in establishing fine-grained, 3D associations between changes in the scene and the agent’s motion planning. Recent advancements in 3D occupancy technologies [15, 16, 17, 18, 19] have gained significant attention from both academia and industry [20, 21]. 3D occupancy offers a well structured 3D representations of surroundings, which facilitates motion planning by establishing fine-grained correlations between surrounding occupied voxels and the agent’s actions.
In this context, the first task is to learn a compact, lower-dimensional representation of the raw occupancy data [4, 3, 5], which must preserve both geometric and semantic details. Current methods mainly employ VQ-VAE [22] for the purpose. While VQ-VAE has shown success in several generative tasks [23, 24, 25], its application to compressing occupancy data forces the latent space of 3D structures to be discretized. This results in different 3D structures sharing the same token in the codebook, inevitably causing a loss of 3D structural features and reduced reconstruction accuracy. Furthermore, in real-world scenarios, most voxels in the occupancy data are empty (unoccupied) [26], and some objects contain only a small number of semantic voxels (e.g., pedestrians, road signs). This leads to the omission of sparse but critical details in the compressed features.
(a) Parallel forecasting.
(b) Autoregressive forecasting.
(c) Performance comparison.
Figure 1: Different principles of occupancy world models: (a) Parallel forecasting [4, 5], which inputs all future queries and historical occupancy results, using a Transformer [27] architecture to predict future occupancy results in parallel; and (b) Autoregressive forecasting [3], which predicts future results iteratively, conditioning each step on a single future query and the previous frame’s prediction. Existing models typically use pre-trained VQ-VAE [22] to compress occupancy results into a compact representative code for predicting future codes. In contrast, our T3Former, shown in (b.2), employs more precise and efficient triplane representations and leverages multi-scale Transformers to predict triplane changes. Therefore, we achieve faster and more accurate occupancy forecasting, as shown in (c).
Based on the scene representation, recent world models [5, 4] often employ Transformers for future state prediction. A key challenge for these approaches is capturing the diverse motion patterns of multi-scale objects on the road. For example, small objects like pedestrians or bicycles may exhibit sudden and abrupt motion changes, while larger objects such as trucks or buses tend to demonstrate more inertial movement, potentially requiring a different modeling granularity. Furthermore, they employ the Transformer to directly predict the complete future occupancy state. This end-to-end approach requires the model to learn both state changes and absolute state reconstruction simultaneously, increasing the learning burden. This not only demands larger model capacity but also increases the risk of error accumulation in long-term predictions.
We present a novel occupancy world model using temporal triplane transformers (T3Former) that addresses the aforementioned challenges. First, we propose to compress the 3D occupancy data into a triplane structure, a concept from computer graphics [28, 29, 30, 31, 32, 33]. A triplane comprises three planes – $xy$, $xz$, and $yz$ – onto which the occupancy data are projected. Compared to VQ-VAE, this representation ensures high reconstruction accuracy of the occupancy grid, achieving a 20% improvement in mIoU while reducing the latent space size by 34%. Next, we extend the representation to the temporal dimension ($xyt$, $xzt$, and $yzt$), using Transformer models across multiple scales to regressively generate incremental changes in the triplane. Finally, we leverage these changes, along with the previous occupancy state, to predict future occupancy and perform motion planning.
Compared to OccWorld, T3Former reduces the total model size by 23%, while also mitigating the cumulative error in long-term occupancy predictions, resulting in a 110.6% increase in mIoU. At the same time, the motion planning accuracy achieves the lowest average error of 1.0 meters and the lowest average collision rate of 30%. This is mainly because we focus on predicting multi-scale occupancy state changes within the triplane space, which more effectively capture environmental dynamics. In summary, our contributions are mainly three-fold:
  • We design a new 4D autoregressive occupancy world model, T3Former, that enhances long-term scene forecasting and enables precise motion planning.
  • T3Former pre-trains a compact occupancy triplane representation to predict future incremental changes of triplanes using multi-scale Transformers. These changes, combined with the previous frame’s output, are then decoded into occupancy states and motion trajectories.
  • Extensive experiments validate our state-of-the-art (SOTA) performance in terms of occupancy forecasting, motion planning, and real-time execution.
2 Related Works
3D occupancy reconstruction. 3D occupancy reconstruction techniques [15, 16, 17, 18, 19, 34] primarily utilze multi-camera RGB images or LiDAR point clouds as inputs to reconstruct the surrounding environment into an agent-centric 3D occupancy space, represented as fixed-size voxels. Each voxel contains a four-tuple attribute [15]: 3D coordinates ($x$, $y$, and $z$) and a semantic category. Existing methods mainly focus on how to establish precise semantic correlations between the raw sensor inputs and the 3D voxels. Once an accurate occupancy reconstruction is achieved, further learning of the temporal dynamics of scene changes becomes necessary [35].
4D occupancy prediction. To infer potential future scene changes, extensive 4D occupancy prediction approaches have been developed to learn the underlying temporal dynamics of scene evolution. Some methods [36, 37, 38] aim to predict future sensor-level data, which is then voxelized into occupancy data, while others [21, 39, 40, 41] leverage historical observations to directly predict occupancy outcomes. These methods mainly focus on reducing spatio-temporal biases on future occupancy predictions. However, they overlook the use of predicted scenes for effective and comprehensive motion planning.
World models for autonomous driving. World models [1, 2] aim to compress high-dimensional scene representations to capture the temporal dynamics of scene transitions, facilitating both future scene predictions and motion planning for the agent. In autonomous driving, existing models [8, 6, 7, 9] typically map surrounding traffic participants to the BEV perspective to predict instance-level tracklets or directly use diffusion models [10, 11, 12, 13, 14] to generate pixel-level future driving views. These methods derive control signals for the agent from current observations and predicted surroundings, but they rely solely on 2D BEV or image space, which limits the ability to establish fine-grained, efficient correlations between scene changes and motion planning. Recent world models [4, 5, 3] have leveraged 3D occupancy data to address this issue. However, they typically use VAE-series [22] models for environment compression, which often neglect original 3D geometric information and compromises reconstruction accuracy. Additionally, they rely on Transformers [27] to forecast the entire future scene instead of incremental changes, leading to significant error accumulation.
3 Methodology
3.1 Formulation
Next, we provide the formulation of Occupancy World Model (OWM) [3]. OWM primarily receives a sequence of scene representations and motion actions from past $\tau_{p}$ frames up to the current timestep $t$, such that $S^{t}\in\mathbb{R}^{H\times W\times L}$ represents the occupancy data of the agent-centric surrounding environment, with $H$, $W$, and $L$ denoting the height, width, and length, respectively, and $a^{t}\in\mathbb{R}^{2}$ denotes a transition-related motion command. The goal of OWM is to establish a stochastic mapping, $\bm{\Phi}$, that associates past occupancy data and actions with future $\tau_{f}$ frames of occupancy data and action proposals. Formally:
$S^{t:t+\tau_{f}},a^{t:t+\tau_{f}}=\bm{\Phi}(S^{t-\tau_{p}:t},a^{t-\tau_{p}:t}).$ (1)
To achieve this, we first pretrain a compressed latent representation of raw occupancy data using an Auto-encoder framework. The encoder and decoder models have parameters $\bm{\Phi}_{enc}$ and $\bm{\Phi}_{\text{dec}}$, respectively. Thus, we have:
$s^{t}=\bm{\Phi}_{\rm enc}(S^{t}),\quad\hat{S}^{t}=\bm{\Phi}_{\rm dec}(s^{t}).$ (2)
where $s^{t}$ and $\hat{S}^{t}$ represent the latent representation and the corresponding reconstructed occupancy data, respectively.
Once we obtain the latent scene representations, we can predict future latent states with incremental changes $\Delta s^{t+1}$. These changes can be aggregated with the data from previous frames and decoded back into occupancy outcomes:
$\Delta s^{t+1},s^{t+1}=\bm{\Phi}_{\rm fut}(s^{t-\tau_{p}:t}),\hat{X}^{t+1}=\bm{\Phi}_{\rm dec}(s^{t+1}).$ (3)
Next, we utilize the incremental latent changes, along with the latent states before the change, and the historical actions, to generate future transitions as follows:
$a^{t+1}=\bm{\Phi}_{\rm act}(a^{t-\tau_{p}:t},s^{t},\Delta s^{t+1}).$ (4)
Note that our OWM, $\bm{\Phi}=\{\bm{\Phi}_{\rm enc},\bm{\Phi}_{\rm dec},\bm{\Phi}_{\rm fut},\bm{\Phi}_{\rm act}\}$, operates in an autoregressive fashion. In particular, Equations (3) and (4) iteratively use previously predicted outcomes as part of the historical data to forecast future scenes and provide motion proposals.
(a) Pretrained traiplane for OWM.
(b) Temporal Triplane Transformers (T3Former) as OWM.
Figure 2: Workflow of T3Former. T3Former first pre-trains compact triplane representations of occupancy data. Next, T3Former utilizes multi-scale Transformers to capture comprehensive temporal dynamics within each plane, thereby predicting future changes in the triplane. Finally, T3Former leverages these predicted changes, along with the previous triplane, to generate future motion proposals.
3.2 T3Former as OWM
3.2.1 Pre-training triplane representations for OWM
To enable faster and more accurate predictions of future scenes $X^{t:t+\tau_{f}}$ by OWM, it is essential to compress the high-dimensional voxelized occupancy scene with high reconstruction accuracy. To achieve this, we employ the Triplane technique [33], which is widely used in volume rendering [28, 29, 30, 31, 32], to compress raw occupancy data in an orthogonal decomposition fashion.
Specifically, as shown in Figure 2(a), the occupancy data $S^{t}$ is first encoded by $\bm{\Phi}_{\rm enc}$ to produce $s^{t}\in\mathbb{R}^{C_{s}\times H_{s}\times W_{s}\times L_{s}}$. Then, an axis-wise average pooling operation is applied to obtain three orthogonal feature planes, $s^{t}=[s^{t}_{xy},s^{t}_{xz},s^{t}_{yz}]$, where $s^{t}_{xy}\in\mathbb{R}^{C_{s}\times W_{s}\times L_{s}}$, $s^{t}_{xz}\in\mathbb{R}^{C_{s}\times H_{s}\times W_{s}}$, and $s^{t}_{yz}\in\mathbb{R}^{C_{s}\times H_{s}\times L_{s}}$, corresponding to the $xy$, $xz$, and $yz$ planes, respectively. To decode the original occupancy data, for each point in the 3D occupancy grid, denoted as $pos=(x,y,z)$, it serves as a position query to retrieve the corresponding features from the three planes in $s^{t}$. These features are then summed and concatenated with the positional encoding ${\rm PE}(pos)$, which is passed into $\bm{\Phi}_{\rm dec}$ to predict the semantic label for the position $pos$. The detailed structures of $\bm{\Phi}_{\rm enc}$ and $\bm{\Phi}_{\rm dec}$ are given in the supplementary martial.
By pretraining $\bm{\Phi}_{\rm enc}$ and $\bm{\Phi}_{\rm dec}$ with Equation (5), we obtain a highly generalizable and accurate triplane representation:
$J_{\rm enc,dec}=\mathbb{E}_{t\sim\mathcal{T},pos\sim S^{t}}[\mathcal{L}_{occ}(\bm{\Phi}_{\rm dec}(\bm{\Phi}_{\rm enc}(S^{t})),S^{t})],$ (5)
where $\mathcal{T}$ represents the collection of all timesteps in the occupancy dataset, and $\mathcal{L}_{occ}=\mathcal{L}_{ce}+\lambda\mathcal{L}_{lz}$. Here, $\mathcal{L}_{ce}$ and $\mathcal{L}_{lz}$ denote the cross-entropy and Lovasz-softmax losses [42, 33], receptively, and $\lambda$ is the trade-off factor.
The triplane representation, compared to the tokens generated by the VQ-VAE in OccWorld [3] and the MS-VAE in OccLLM [5], retains 3D structural information while achieving a more compact latent space. This not only makes our OWM more lightweight but also reduces the cumulative prediction error over time, as detailed later.
3.2.2 Multi-scale autoregressions in T3Former
Our T3Former model functions similarly to GPT-series [43, 44, 45], leveraging historical triplanes to generate future triplanes and using each predicted triplane to iteratively forecast subsequent ones. At timestep $t$, given the historical $\tau_{p}$ frames of triplanes $s^{t-\tau_{p}:t}$, the goal of T3Former is to capture the full temporal dynamics of the scene, particularly adapting to the motion patterns of objects of varying sizes. To achieve this, T3Former follows two key steps: (i) predicting each plane’s future changes, $\{\Delta s^{t:t+\tau_{f}}_{i}\mid i\in\{xy,xz,yz\}\}$, using Transformers with multiple scales; and (ii) aggregating these changes into the previous plane state and aligning the three plane predictions by fine-tuning the decoder $\bm{\Phi}_{\rm dec}$.
Figure 3: Motion planning with T3Former.
In particular, $\bm{\Phi}_{\rm fut}=\{\bm{\Phi}_{{\rm fut}_{xy}},\bm{\Phi}_{{\rm fut}_{xz}},\bm{\Phi}_{{\rm fut}_{yz}}\}$. Each plane’s prediction model, $\bm{\Phi}_{{\rm fut}_{i}}$, consists of Transformers [27] operating at multiple scales. These Transformers share the same architecture but have distinct learnable parameters and varying input sizes, depending on the plane and scale, as shown in Figure 2(b). For example, in the case of $\bm{\Phi}_{{\rm fut}_{xy}}$, the input $s^{t}_{xy}$ is passed through UNet [46]-style downsampling, resulting in $V$ scales of plane features. The feature at each scale $v\in V$ is denoted as $s^{t,v}_{xy}\in\mathbb{R}^{C_{s_{xy}}^{v}\times W_{s_{xy}}^{v}\times H_{s_{xy}}^{v}}$. These features are then flattened across the last two dimensions to form $W_{s_{xy}}^{v}\times H_{s_{xy}}^{v}$ tokens, which are input into the Transformer encoder to generate spatio-temporal memory. Next, we employ a learnable query $Q^{t,v}_{s_{xy}}$, which has the same dimension as $s^{t,v}_{xy}$, to enable cross-attention in the Transformer decoder and generate the token change from the current to the next time step based on the memory. Finally, the token changes at all scales are aggregated using UNet-style upsampling, yielding the plane’s feature change at the next time step, $\Delta\hat{s}^{t+1}_{xy}$. This change is then combined with the previous plane state, $\hat{s}^{t}_{xy}$, through a convolutional operation $\phi$ with a $1\times 1$ kernel size.
The autoregressive forecasting process is given by:
$\Delta\hat{s}^{t+k}_{i}=\bm{\Phi}_{{\rm fut}_{i}}(\hat{s}^{t+k-\tau_{p}:t+k}_{i},Q^{t+k}_{s_{i}}),$ (6)
$\hat{s}^{t+k}_{i}=\Delta\hat{s}^{t+k}_{i}+\phi(\hat{s}^{t+k-1}_{i}),$ (7)
$\hat{S}^{t+k}=\bm{\Phi}_{\rm dec}(\hat{s}^{t+k}:=\{\hat{s}^{t+k}_{i}\}),$ (8)
where $i\in\{xy,xz,yz\}$, $k\in\{1,\cdots,\tau_{f}\}$ is the timestep of autoregressive forecasting, and $Q^{t+k}_{s_{i}}$ is the learnable query at timestep $t+k$ in the $i$-th plane.
Since we have pre-trained the triplane occupancy representations, the future triplane can be used as the ground truth (GT) to supervise T3Former. Besides, we finetune $\Phi_{\rm dec}$ to the separate predictions of each plane. Formally:
$J_{\rm fut}=\mathbb{E}_{t,k,i}[\mathcal{L}_{\rm fut}(\hat{s}^{t+k}_{i},s^{t+k}_{i})+\xi\mathcal{L}_{\rm occ}(\hat{S}^{t+k},S^{t+k})],$ (9)
where $\mathcal{L}_{\rm fut}$ is the weighted sum of L1 and L2 losses, defined as $\mathcal{L}_{\rm fut}=\mathcal{L}_{1}+\gamma\mathcal{L}_{2}$, and $\gamma,\xi$ are trade-off weights.
3.3 Motion planning with T3Former
Given the predicted scene changes, $\Delta\hat{s}^{t+k}$, provided by T3Former, we can generate a comprehensive motion plan by anticipating potential dangers based on both the scene changes and the previous and next scenes.
To achieve this, we first map the previous frame triplane $\hat{s}^{t+k-1}$, the triplane change $\Delta\hat{s}^{t+k}$, and the next frame triplane $\hat{s}^{t+k}$ through the ResNet-18 networks [47] $\theta_{p}$, $\theta_{\Delta}$, and $\theta_{f}$, respectively, into a shared latent space, denoted as $z^{t+k-1}$, $\Delta z^{t+k}$, and $z^{t+k}$. Then, $\Delta z^{t+k}$ passes through two independent fully connected (FC) layers and Sigmoid layers, collectively referred to as $\zeta_{p}$ and $\zeta_{f}$, to generate the query vectors $f^{t+k-1}$ and $f^{t+k}$, which are then multiplied with $z^{t+k-1}$ and $z^{t+k}$ to obtain the motion-related features of the scene change relative to the last triplane and the next triplane. Finally, the motion-related features and the scene change $\Delta z^{t+k}$ are element-wise averaged and processed through a FC layer, and then added to the future position encoding, resulting in the query variable for the next timestep motion plan, $Q^{t+k}_{a}\in\mathbb{R}^{d_{\rm act}}$.
Next, we project the actions of the historical $\tau_{p}$ frames into the same dimension as $Q^{t+k}_{a}$, and employ a Transformer encoder to capture the motion dependencies within these frames. Then, $Q^{t+k}_{a}$ is used in the Transformer decoder via cross-attention to predict the next action $\hat{a}^{t+k+1}$. The autoregressive generation of motion plans is defined as:
$\hat{a}^{t+k+1}=\bm{\Phi}_{{\rm act}}(\hat{a}^{t+k-\tau_{p}:t+k},\hat{s}^{t+k-1},\Delta\hat{s}^{t+k},Q^{t+k}_{a}).$ (10)
Unlike OccWorld [3], we do not require an additional ego token to continuously track the agent’s motion trajectory. Instead, we can learn future motion directly from the scene changes and historical motion, as shown in Figure 3.
The optimization objective is defined as follows:
$J_{\rm act}=\mathbb{E}_{t,k}[\mathcal{L}_{\rm act}(\hat{a}^{t+k},a^{t+k})],$ (11)
where $\mathcal{L}_{\rm act}$ measures the L2 discrepancy between the predicted and GT trajectories.
Models Input mIoU (%) $\uparrow$ IoU (%) $\uparrow$
0s 1s 2s 3s Avg. 0s 1s 2s 3s Avg.
Copy&Paste 3D-Occ 66.38 14.91 10.54 8.52 11.33 62.29 24.47 19.77 17.31 20.52
OccWorld-O [3] 3D-Occ 66.38 25.78 15.14 10.51 17.14 62.29 34.63 25.07 20.18 26.63
Vanilla-LLaMA-O [4] 3D-Occ - 14.15 9.80 6.77 10.24 - 21.36 18.31 14.82 18.16
OccLLaMA-O [4] 3D-Occ 75.20 25.05 19.49 15.26 19.93 63.76 34.56 28.53 24.41 29.17
OccLLM-O [5] 3D-Occ - 24.02 21.65 17.29 20.99 - 36.65 32.14 28.77 32.52
T3Former-O 3D-Occ 85.50 46.32 33.23 28.73 36.09 92.07 77.00 75.89 76.32 76.40
OccWorld-F [3, 48] Camera 20.09 8.03 6.91 3.54 6.16 35.61 23.62 18.13 15.22 18.99
OccLLaMA-F [4, 48] Camera 37.38 10.34 8.66 6.98 8.66 38.92 25.81 23.19 19.97 22.99
OccLLM-F [5, 48] Camera - 11.28 10.21 9.13 10.21 - 27.11 24.07 20.19 23.79
T3Former-F [48] Camera 43.52 24.87 18.30 15.63 19.60 54.31 38.98 37.45 31.89 36.11
Table 1: Testing performance comparison with SOTA methods on the 4D occupancy forecasting task. Best values in each metric are bolded. 0s refers to reconstruction accuracy, while 1s, 2s, and 3s denote future prediction accuracy. Avg. is the average of 1s, 2s, and 3s.
4 Experiments
4.1 Experimental settings
Targets and evaluation metrics. The main focus of OWM is jointly modeling occupancy forecasting and motion planning. Following the conventions in [3, 5, 4], we use the past four frames (2 seconds) to predict the outcomes of the next six frames (3 seconds). We conduct two sets of experiments:
  • To validate the accuracy of 4D occupancy forecasting, we use intersection over union (IoU) and mean IoU (mIoU) across all semantic classes to measure the predicted future occupancy under the Occ3D dataset [17].
  • To assess the precision and safety of motion planning, we calculate the L2 distance (in meters) between the planned and GT trajectories, and the collision rate with traffic participants’ bounding boxes using the nuScenes dataset [49].
Models Input Auxiliary supervision L2 (m) $\downarrow$ Collision rate (%) $\downarrow$
1s 2s 3s Avg. 1s 2s 3s Avg.
IL [50] LiDAR None 0.44 1.15 2.47 1.35 0.08 0.27 1.95 0.77
NMP [51] LiDAR Box+Motion 0.53 1.25 2.67 1.48 0.04 0.12 0.87 0.34
FF [52] LiDAR Freespace 0.55 1.20 2.54 1.43 0.06 0.17 1.07 0.43
EO [53] LiDAR Freespace 0.67 1.36 2.78 1.60 0.04 0.09 0.88 0.33
ST-P3 [8] Camera Map+Box+Depth 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71
UniAD [6] Camera Map+Box+Motion+Track+Occ 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31
VAD [7] Camera Map+Box+Motion 0.54 1.15 1.98 1.22 0.04 0.39 1.17 0.53
OccNet [15] Camera Map+Box+3D-Occ 1.29 2.13 2.99 2.14 0.21 0.59 1.37 0.72
OccNet [15] 3D-Occ Map+Box 1.29 2.31 2.98 2.25 0.20 0.56 1.30 0.69
OccWorld-O [3] 3D-Occ None 0.43 1.08 1.99 1.17 0.07 0.38 1.35 0.60
OccLLaMA-O [4] 3D-Occ None 0.37 1.02 2.03 1.14 0.04 0.24 1.20 0.49
T3Former-O 3D-Occ None 0.32 0.91 1.76 1.00 0.08 0.32 0.51 0.30
Table 2: Testing performance of motion planning compared with SOTA method. Best and second-best values in each metric are bolded and underlined, respectively. Auxiliary supervision refers to additional supervision signals beyond the GT trajectories.
Implementation details. The dataset consists of 1,000 scenes, of which 850 are used for training and 100 for testing. Each scene contains up to 40 timesteps, with a sampling frequency of 2Hz. The dimensions of the occupancy data $S^{t}$ at each timestep are $16\times 200\times 200$. The pre-trained Triplane feature dimensions are $8\times 16\times 100\times 100$. In $\bm{\Phi}_{\rm fut}$, predictions for each plane incorporate features at $V=5$ different scales, while the token dimension in $\bm{\Phi}_{\rm act}$ is $d_{\rm act}=50$. The objectives, $J_{\rm enc,dec}$, $J_{\rm fut}$, and $J_{\rm mot}$, are optimized using AdamW with a weight regularization factor of 0.01, an initial learning rate of 0.001, and cosine decay with a minimum learning rate of $10^{-6}$. We first pre-train $\bm{\Phi}_{\rm enc}$ and $\Phi_{\rm dec}$ with a batch size of 10, using random flip augmentation to obtain the triplane representations. Then, we train $\Phi_{\rm fut}$ and $\bm{\Phi}_{\rm act}$ with a batch size of 1, while fine-tuning $\bm{\Phi}_{\rm dec}$. All training and testing are performed on 4 RTX 4090 GPUs. More details are provided in the supplementary material.
4.2 Comparisons with the state-of-the-art
4.2.1 4D occupancy forecasting
Table 1 presents the performance results of various methods, considering two modes of operation: (i) using 3D occupancy GTs as historical input, marked with “-O”; (ii) using predicted 3D occupancy data from FB-Occ [48] as historical input, marked with “-F”. “Copy&Paste” refers to directly using the GT occupancy data from the current timestep as predictions for future outcomes. Vanilla-LLaMA-O refers to the method that simply flattens the occupancy data and feeds it into LLaMA [54] for training.
It is clear that SOTA methods’ $\bm{\Phi}_{\rm enc}$ fail to achieve high-fidelity compression, resulting in low reconstruction accuracy at time 0s. In contrast, both T3Former-O and T3Former-F utilize triplane representations that preserve both geometric and semantic information, leading to the highest mIoU and IoU reconstruction accuracy. This significantly aids in $\bm{\Phi}_{\rm fut}$ for predicting the next 3 seconds, effectively reducing error accumulation. Specifically, the “Copy&Paste” method, as a simple baseline, demonstrates substantial error accumulation. In contrast, T3Former-O shows a significant reduction in error, as indicated by mIoU and IoU averages. Notably, IoU generally surpasses mIoU, as predicting occupancy (whether occupied or unoccupied) is relatively straightforward, while semantic tracking without instance-level supervision is more challenging and prone to drift over time.
4.2.2 Motion planning
We extensively compare T3Former with SOTA methods for autonomous driving, including LiDAR-based methods (IL [50], NMP [51], FF [52], and EO [53]), camera-based methods (ST-P3 [8], UniAD [6], VAD [7], and OccNet [15]), and occupancy-based methods (OccNet [15], OccWorld [3], and OccLLaMA [4]). LiDAR- and camera-based methods explore various combinations of auxiliary supervisions to enhance the quality of planned motion trajectories. Despite their success, the auxiliary information they rely on is labor-intensive to obtain.
In contrast, occupancy-based methods generate motion proposals using only occupancy representations. T3Former achieves the lowest L2 error over time, the lowest collision rate at the final time step, and the lowest average collision rate across all future motions. Compared to SOTA occupancy-based methods, T3Former reduces the average L2 error by 12.3% and the average collision rate by 38.8%. This is primarily because T3Former can establish precise, task-relevant correlations between scene changes and motion trajectories in a deep latent space, effectively filtering out noise from scene variations.
Figure 4: Visualization of 4D occupancy forecasting and motion planning for the next 3 seconds (zoom in for a clearer view).
4.3 Visualizations
Comparisons with SOTA. Figure 4 illustrates the evolution of the scene over the next 3 seconds, along with the corresponding motion predictions for two different scenarios. (i) In the first case, it is clear that OccWorld-O incorrectly labels the truck as a trailer initially, resulting in a continuous accumulation of errors. In contrast, T3Former effectively learns the motion dynamics of all objects in the scene, consistently tracking each vehicle over a long horizon, while providing clearer object boundaries. (ii) In the second case, as the prediction horizon extends, OccWorld-O’s road boundary gradually fades, losing its clarity, whereas our method maintains the boundary shape of the road, including the trees. Additionally, OccWorld-O misidentifies the motorcycle as a truck at both 2.5s and 3s, while T3Former-O ensures that the semantics of small objects remain accurate over time. (iii) Under both scenarios, T3Former-O enables precise long-term future predictions, allowing our motion planning to rely solely on the evolving scene dynamics. This leads to more accurate trajectories and better avoidance of potential future hazards.
Driving view generation. Recent diffusion-based autonomous driving models [12, 13, 11] typically use historical driving views as conditions to generate future views. In contrast, our approach uses only triplane data as a lightweight condition. We fine-tune the diffusion model with ControlNet [55]. Figure 5 showcases synthetic images generated with different weather prompts and different triplane conditions. Triplane provides accurate 3D structural information, thereby enhancing consistency between pixel positions and semantics in the resulting synthetic 2D images.
Figure 5: Driving view generations conditioned on different weather prompts and different triplane constraints.
Figure 6: Predicted locations of dynamic objects for the three last frames in the prediction window.
Figure 7: A failure case at a T-junction scenario.
Comparison of motion prediction. Figure 6 compares occupancy predictions for the three furthest frames in the prediction window, focusing on dynamic objects of varying sizes (cars and pedestrians) from a BEV perspective. Two reference lines are displayed to reflect the errors between the predictions and GT. The results of our method match the GT significantly better than OccWorld-O, which exhibits substantial drift. For instance, the predictions for the pedestrians at the top and the vehicle on the left clearly demonstrate this improvement
Failure case. Figure 7 shows a failure case across three frames, where the road boundary at the end contains sparse occupied voxels, making the prediction challenging. Due to the inherent blurring issue of VAE [56, 57], OccWorld predictions tend to smooth and homogenize the sparse regions. In contrast, the triplane consists of fine-grained geometric details, which causes our predictions to also exhibit sparse boundaries, resulting in some deviation in the trajectory.
4.4 Ablation study
To validate the effectiveness of the key designs in T3Former, we conduct ablation studies by evaluating the averaged mIoU and averaged L2 across future outcomes, as shown in Table 3. M0 represents our complete version. For fairness, all ablation models are derived by replacing specific modules in M0, as outlined below: M1 replaces the input of T3Former with the VQ-VAE token from OccWorld [3], which results in a significant performance drop. M2 uses a single-scale Transformer to learn temporal dynamics based on triplane representations, which fails to capture the motion patterns of multi-scale objects, leading to poor performance. M3 leverages multi-scale Transformers for parallel forecasting, as shown in Figure 1(a), instead of autoregression, which leads to error accumulation. M4 directly predicts the entire future triplane, demonstrating that predicting changes in the triplane is more efficient.
Idx. Models Avg. mIoU $\uparrow$ Avg. L2 $\downarrow$
M0 T3Former-O 34.70 0.63
M1 w/o Triplane 25.76 0.96
M2 w/o Multi-scale mot. 31.92 0.87
M3 w/o Autoregression 32.44 0.79
M4 w/o Changing triplane 33.81 0.71
Table 3: Ablation study of the components in T3Former.
4.5 Efficiency analysis
The predefined latent space shapes of OccWorld [3] and OccLLM [5] are 128$\times$50$\times$50 and 64$\times$50$\times$50, respectively. In contrast, T3Former only requires a compact triplane representation that is in spirit similar to orthogonal decomposition [33, 32]. Specifically, the $xy$ plane has a shape of 8$\times$100$\times$100, while the $xz$ and $yz$ planes are represented with shapes of 8$\times$100$\times$16. This results in the smallest latent shape and parameter count for $\Phi_{\rm enc,dec}$, while achieving superior reconstruction accuracy, as shown in Table 4.
Models Latent shape $\downarrow$ Params (M) $\downarrow$ Reconstruction
mIoU $\uparrow$ IoU $\uparrow$
OccWorld [3] 320,000 14.17 60.50 59.07
OccLLM [5] 160,000 2.30 71.08 62.74
T3Former 105,600 0.69 85.50 92.07
Table 4: Comparisons of latent representations within OWMs.
Based on the triplane representation learned by $\Phi_{\rm enc}$, our $\Phi_{\rm fut}$ only needs to learn the changes in the triplane, while $\Phi_{\rm act}$ predicts the motion trajectory based on these changes. This results in our lightweight design, with a total parameter count of 55M, compared to OccWorld’s 72M, achieving a 23% reduction. As shown in Figure 1(c), T3Former achieves 26 FPS for predicting each future frame on an RTX 4090 GPU, whereas OccWorld delivers only 18 FPS.
5 Conclusion
This paper proposes a new 4D occupancy world model, T3Former, which features three key designs: (i) pre-training triplane representations for OWM, (ii) capturing comprehensive motion dynamics using multi-scale Transformers, and (iii) predicting triplane changes for occupancy forecasting and motion planning. T3Former achieves SOTA performance in 4D occupancy forecasting and motion planning, with real-time capability. Our future work will dive deeper into the generation of driving views, as shown in Figure 5, focusing on the synergy between prompt constraints, image texture, and T3Former.