### v2

# Temporal Triplane Transformers as Occupancy World Models

Abstract
World models aim to learn or construct representations of the environment that enable the prediction of future scenes, thereby supporting intelligent motion planning. However, existing models often struggle to produce fine-grained predictions and to operate in real time. In this work, we propose T3Former, a novel 4D occupancy world model for autonomous driving. T3Former begins by pre-training a compact triplane representation that efficiently encodes 3D occupancy. It then extracts multi-scale temporal motion features from historical triplanes and employs an autoregressive approach to iteratively predict future triplane changes. Finally, these triplane changes are combined with previous states to decode future occupancy and ego-motion trajectories. Experimental results show that T3Former achieves 1.44$\times$ speedup (26 FPS), improves mean IoU to 36.09, and reduces mean absolute planning error to 1.0 meters. Demos are available in the supplementary material.
1 Introduction
World models [10, 9] aim to represent the environment, predict future scenes and enable agents to perform advanced motion planning. These models first encode sensory inputs into compact, lower-dimensional representations, which are then used to model the temporal evolution of the environment. When applied to autonomous driving [58, 46, 51, 16, 8, 53], world models help anticipate potential hazards and support more informed, safer decision-making [6].
Existing models [14, 16, 19, 57] focus on predicting future instance segmentation of traffic participants from a Bird’s Eye View (BEV), or use diffusion models [48, 45, 44, 7, 18] to generate future driving scenes. However, these approaches encounter challenges in establishing fine-grained, 3D associations between scene changes and the agent’s motion planning. Recently, 3D occupancy  [40, 32, 39, 47, 28] has emerged as a structured and spatially rich representation of the environment. It enables finer correlations between occupied voxels and planning decisions, thus offering improved support for downstream control tasks. In this paper, we focus on the design of Occupancy World Models (OWMs) to address this challenge.
(a) Conventional approach.
(b) Our approach T3Former.
Figure 1: Different occupancy world architectures: Conventional approaches [58, 46, 51, 53, 8], as shown in (a), use a pre-trained VAE [41, 24] to compress occupancy data (denoted by $S$) into a compact representation code, which is then combined with encoded previous actions (marked as $a$) to predict future codes and actions through the Transformer. In contrast, our method, T3Former, shown in (b), employs more precise and efficient triplane representations and leverages a lightweight Transformer model to separately predict the triplane changes of each plane ($\Delta$). These $\Delta$ values are then used to for occupancy forecasting and motion planning.
In this context, the first task is to learn a compact representation of the raw occupancy data [46, 58, 51], which must preserve both geometric and semantic details. Current methods mainly employ VQ-VAE [41, 24] for the purpose. While VQ-VAE has shown success in several generative tasks [35, 20, 55], its application to compressing occupancy data forces the latent space of 3D structures to be discretized. This results in different 3D structures sharing the same token in the codebook, inevitably causing a loss of 3D structural features and reduced accuracy. Furthermore, in real-world scenarios, most voxels in the occupancy data are empty (unoccupied) [43], and some objects contain only a small number of semantic voxels (e.g., pedestrians, road signs). This leads to the omission of sparse but critical details in the compressed features.
Figure 2: Performance comparisons on the test dataset. Our T3Former achieves more accurate occupancy forecasting and motion planning with faster inference speed.
Based on the scene representation, recent world models [51, 46] often use Transformers for future state prediction. A key challenge here is capturing the diverse motion patterns of multi-scale objects on the road. For example, small objects like pedestrians or bicycles may exhibit sudden and abrupt motion changes, while larger objects such as trucks or buses tend to demonstrate more inertial movement, potentially requiring a different modeling granularity. Furthermore, they employ the Transformer to directly predict the complete future occupancy state. This end-to-end approach requires the model to learn both state changes and absolute state reconstruction simultaneously, increasing the learning burden. This not only demands larger model capacity but also increases the risk of error accumulation in long-term predictions.
We present a novel occupancy world model using temporal triplane transformers, termed T3Former, that addresses the aforementioned challenges. First, we propose to compress the 3D occupancy data into a triplane structure, a concept from computer graphics [21, 37, 15, 38, 17, 25]. A triplane comprises three planes – $xy$, $xz$, and $yz$ – onto which the occupancy data are projected. Compared to VQ-VAE, this representation ensures high reconstruction accuracy of the occupancy grid, achieving a 20% improvement in mIoU while reducing the latent space size by 34%. Next, we extend the representation to the temporal dimension ($xyt$, $xzt$, and $yzt$), using Transformer models across multiple scales to regressively generate incremental changes in the triplane. Finally, we leverage these changes, along with the previous occupancy state, to predict future occupancy and perform motion planning.
Compared to conventional approaches, we focus on predicting occupancy state changes across varying scales within the triplane space, which more effectively capture environmental dynamics. As a result, as shown in Figure 2, T3Former mitigates cumulative errors in long-term occupancy predictions, leading to an 81.1% increase in averaged mIoU. Additionally, motion planning accuracy achieves the lowest average error of 1.0 meter and the lowest average collision rate of 30%. Furthermore, by leveraging the concept of $\Delta$ predictions, T3Former significantly reduces model capacity, achieving 26 FPS. In summary, our contributions are mainly three-fold:
$\bullet$ We design a novel 4D autoregressive occupancy world model, T3Former, that enhances long-term scene forecasting and enables precise motion planning.
$\bullet$ T3Former pre-trains a compact occupancy triplane representation to predict future incremental changes of triplanes using multi-scale Transformers. These changes, combined with the previous frame’s output, are then decoded into occupancy states and motion trajectories.
$\bullet$ Extensive experiments validate our state-of-the-art (SOTA) performance in terms of occupancy forecasting, motion planning, and real-time execution.
2 Related Works
3D occupancy reconstruction. 3D occupancy reconstruction techniques [40, 32, 39, 47, 28, 52] primarily utilze multi-camera RGB images or LiDAR point clouds as inputs to reconstruct the surrounding environment into an agent-centric 3D occupancy space, represented as fixed-size voxels. Each voxel contains a four-tuple attribute [40]: 3D coordinates ($x$, $y$, and $z$) and a semantic category. Existing methods mainly focus on how to establish precise semantic correlations between the raw sensor inputs and the 3D voxels. Once an accurate occupancy reconstruction is achieved, further learning of the temporal dynamics of scene changes becomes necessary [26].
4D occupancy prediction. To infer potential future scene changes, extensive 4D occupancy prediction approaches have been developed to learn the underlying temporal dynamics of scene evolution. Some methods [30, 33, 23] aim to predict future sensor-level data, which is then voxelized into occupancy data, while others [31, 4, 50, 49] leverage historical observations to directly predict occupancy outcomes. These methods mainly focus on reducing spatio-temporal biases on future occupancy predictions. However, they overlook the use of predicted scenes for effective and comprehensive motion planning.
World models for autonomous driving. World models [10, 9] aim to compress high-dimensional scene representations to capture the temporal dynamics of scene transitions, facilitating both future scene predictions and motion planning for the agent. In autonomous driving, existing models [14, 16, 19, 57] typically map surrounding traffic participants to the BEV perspective to predict instance-level tracklets or directly use diffusion models [48, 45, 44, 7, 18] to generate pixel-level future driving views. These methods derive control signals for the agent from current observations and predicted surroundings, but they rely solely on 2D BEV or image space, which limits the ability to establish fine-grained, efficient correlations between scene changes and motion planning. Recent world models [46, 51, 58] have leveraged 3D occupancy data to address this issue. However, they typically use VAE-series [41] models for environment compression, which often neglect original 3D geometric information and compromises reconstruction accuracy. Additionally, they rely on Transformers [42] to forecast the entire future scene instead of incremental changes, leading to significant error accumulation.
3 Methodology
3.1 Formulation
Next, we provide the formulation of Occupancy World Model (OWM) [58]. OWM primarily receives a sequence of scene representations and motion actions from past $\tau_{p}$ frames up to the current timestep $t$, such that $S^{t}\in\mathbb{R}^{H\times W\times L}$ represents the occupancy data of the agent-centric surrounding environment, with $H$, $W$, and $L$ denoting the height, width, and length, respectively, and $a^{t}\in\mathbb{R}^{2}$ denotes a transition-related motion command. The goal of OWM is to establish a stochastic mapping, $\bm{\Phi}$, that associates past occupancy data and actions with future $\tau_{f}$ frames of occupancy data and action proposals. Formally:
$S^{t:t+\tau_{f}},a^{t:t+\tau_{f}}=\bm{\Phi}(S^{t-\tau_{p}:t},a^{t-\tau_{p}:t}).$ (1)
To achieve this, we first pretrain a compressed latent representation of raw occupancy data using an Auto-encoder framework. The encoder and decoder models have parameters $\bm{\Phi}_{enc}$ and $\bm{\Phi}_{\text{dec}}$, respectively. Thus, we have:
$s^{t}=\bm{\Phi}_{\rm enc}(S^{t}),\quad\hat{S}^{t}=\bm{\Phi}_{\rm dec}(s^{t}).$ (2)
where $s^{t}$ and $\hat{S}^{t}$ represent the latent representation and the corresponding reconstructed occupancy data, respectively.
Once we obtain the latent scene representations, we can predict future latent states with incremental changes $\Delta s^{t+1}$. These changes can be aggregated with the data from previous frames and decoded back into occupancy outcomes:
$\Delta s^{t+1},s^{t+1}=\bm{\Phi}_{\rm fut}(s^{t-\tau_{p}:t}),\hat{X}^{t+1}=\bm{\Phi}_{\rm dec}(s^{t+1}).$ (3)
Next, we utilize the incremental latent changes, along with the latent states before the change, and the historical actions, to generate future transitions as follows:
$a^{t+1}=\bm{\Phi}_{\rm act}(a^{t-\tau_{p}:t},s^{t},\Delta s^{t+1}).$ (4)
Note that our OWM, $\bm{\Phi}=\{\bm{\Phi}_{\rm enc},\bm{\Phi}_{\rm dec},\bm{\Phi}_{\rm fut},\bm{\Phi}_{\rm act}\}$, operates in an autoregressive fashion. In particular, Eqs. (3) and (4) iteratively use previously predicted outcomes as part of the historical data to forecast future scenes and provide motion proposals. Figure 3 illustrates our design, detailed below.
Figure 3: Workflow of T3Former. T3Former first pre-trains compact triplane representations of occupancy data. Next, T3Former utilizes multi-scale Transformers to capture comprehensive temporal dynamics within each plane, thereby predicting future triplane changes. Finally, T3Former leverages these predicted changes, along with the previous triplane, to generate future motion proposals.
3.2 T3Former as OWM
3.2.1 Pre-training triplane representations for OWM
Figure 4: Pretrained traiplane for OWM.
To enable faster and more accurate predictions of future scenes $X^{t:t+\tau_{f}}$ by OWM, it is essential to compress the high-dimensional voxelized occupancy scene with high reconstruction accuracy. To achieve this, we employ the Triplane technique [25], which is widely used in volume rendering [21, 37, 15, 38, 17], to compress raw occupancy data in an orthogonal decomposition fashion.
Specifically, as shown in Figure 4, the occupancy data $S^{t}$ is first encoded by $\bm{\Phi}_{\rm enc}$ to produce $s^{t}\in\mathbb{R}^{C_{s}\times H_{s}\times W_{s}\times L_{s}}$. Then, an axis-wise average pooling operation is applied to obtain three orthogonal feature planes, $s^{t}=[s^{t}_{xy},s^{t}_{xz},s^{t}_{yz}]$, where $s^{t}_{xy}\in\mathbb{R}^{C_{s}\times W_{s}\times L_{s}}$, $s^{t}_{xz}\in\mathbb{R}^{C_{s}\times H_{s}\times W_{s}}$, and $s^{t}_{yz}\in\mathbb{R}^{C_{s}\times H_{s}\times L_{s}}$, corresponding to the $xy$, $xz$, and $yz$ planes, respectively. To decode the original occupancy data, for each point in the 3D occupancy grid, denoted as $pos=(x,y,z)$, it serves as a position query to retrieve the corresponding features from the three planes in $s^{t}$. These features are then summed and concatenated with the positional encoding ${\rm PE}(pos)$, which is passed into $\bm{\Phi}_{\rm dec}$ to predict the semantic label for the position $pos$. The detailed structures of $\bm{\Phi}_{\rm enc}$ and $\bm{\Phi}_{\rm dec}$ are given in the supplementary martial.
By pretraining $\bm{\Phi}_{\rm enc}$ and $\bm{\Phi}_{\rm dec}$ with Eq. (5), we obtain a highly generalizable and accurate triplane representation:
$J_{\rm enc,dec}=\mathbb{E}_{t\sim\mathcal{T},pos\sim S^{t}}[\mathcal{L}_{occ}(\bm{\Phi}_{\rm dec}(\bm{\Phi}_{\rm enc}(S^{t})),S^{t})],$ (5)
where $\mathcal{T}$ represents the collection of all timesteps in the occupancy dataset, and $\mathcal{L}_{occ}=\mathcal{L}_{ce}+\lambda\mathcal{L}_{lz}$. Here, $\mathcal{L}_{ce}$ and $\mathcal{L}_{lz}$ denote the cross-entropy and Lovasz-softmax losses [1, 25], receptively, and $\lambda$ is the trade-off factor.
The triplane representation, compared to the tokens generated by the VQ-VAE in OccWorld [58] and the MS-VAE in OccLLM [51], retains 3D structural information while achieving a more compact latent space. This not only makes our OWM more lightweight but also reduces the cumulative prediction error over time, as detailed later.
3.2.2 Multi-scale autoregressions in T3Former
Our T3Former model functions similarly to GPT-series [2, 29, 59], leveraging historical triplanes to generate future triplanes and using each predicted triplane to iteratively forecast subsequent ones. At timestep $t$, given the historical $\tau_{p}$ frames of triplanes $s^{t-\tau_{p}:t}$, the goal of T3Former is to capture the full temporal dynamics of the scene, particularly adapting to the motion patterns of objects of varying sizes. To achieve this, T3Former follows two key steps: (i) predicting each plane’s future changes, $\{\Delta s^{t:t+\tau_{f}}_{i}\mid i\in\{xy,xz,yz\}\}$, using Transformers with multiple scales; and (ii) aggregating these changes into the previous plane state and aligning the three plane predictions by fine-tuning the decoder $\bm{\Phi}_{\rm dec}$.
In particular, $\bm{\Phi}_{\rm fut}=\{\bm{\Phi}_{{\rm fut}_{xy}},\bm{\Phi}_{{\rm fut}_{xz}},\bm{\Phi}_{{\rm fut}_{yz}}\}$. Each plane’s prediction model, $\bm{\Phi}_{{\rm fut}_{i}}$, consists of Transformers [42] operating at multiple scales. These Transformers share the same architecture but have distinct learnable parameters and varying input sizes, depending on the plane and scale, as shown in Figure 3. For example, in the case of $\bm{\Phi}_{{\rm fut}_{xy}}$, the input $s^{t}_{xy}$ is passed through UNet [36]-style downsampling, resulting in $V$ scales of plane features. The feature at each scale $v\in V$ is denoted as $s^{t,v}_{xy}\in\mathbb{R}^{C_{s_{xy}}^{v}\times W_{s_{xy}}^{v}\times H_{s_{xy}}^{v}}$. These features are then flattened across the last two dimensions to form $W_{s_{xy}}^{v}\times H_{s_{xy}}^{v}$ tokens, which are input into the Transformer encoder to generate spatio-temporal memory. Next, we employ a learnable query $Q^{t,v}_{s_{xy}}$, which has the same dimension as $s^{t,v}_{xy}$, to enable cross-attention in the Transformer decoder and generate the token change from the current to the next time step based on the memory. Finally, the token changes at all scales are aggregated using UNet-style upsampling, yielding the plane’s feature change at the next time step, $\Delta\hat{s}^{t+1}_{xy}$. This change is then combined with the previous plane state, $\hat{s}^{t}_{xy}$, through a convolutional operation $\phi$ with a $1\times 1$ kernel size.
The autoregressive forecasting process is given by:
$\Delta\hat{s}^{t+k}_{i}=\bm{\Phi}_{{\rm fut}_{i}}(\hat{s}^{t+k-\tau_{p}:t+k}_{i},Q^{t+k}_{s_{i}}),$ (6)
$\hat{s}^{t+k}_{i}=\Delta\hat{s}^{t+k}_{i}+\phi(\hat{s}^{t+k-1}_{i}),$ (7)
$\hat{S}^{t+k}=\bm{\Phi}_{\rm dec}(\hat{s}^{t+k}:=\{\hat{s}^{t+k}_{i}\}),$ (8)
where $i\in\{xy,xz,yz\}$, $k\in\{1,\cdots,\tau_{f}\}$ is the timestep of autoregressive forecasting, and $Q^{t+k}_{s_{i}}$ is the learnable query at timestep $t+k$ in the $i$-th plane.
Since we have pre-trained the triplane occupancy representations, the future triplane can be used as the ground truth (GT) to supervise T3Former. Additionally, as we predict each plane independently, misalignments may occur. We finetune $\Phi_{\rm dec}$ to the align predictions of each plane. Formally:
$J_{\rm fut}=\mathbb{E}_{t,k,i}[\mathcal{L}_{\rm fut}(\hat{s}^{t+k}_{i},s^{t+k}_{i})+\xi\mathcal{L}_{\rm occ}(\hat{S}^{t+k},S^{t+k})],$ (9)
where $\mathcal{L}_{\rm fut}$ is the weighted sum of L1 and L2 losses, defined as $\mathcal{L}_{\rm fut}=\mathcal{L}_{1}+\gamma\mathcal{L}_{2}$, and $\gamma,\xi$ are trade-off weights.
3.3 Motion planning with T3Former
Figure 5: Motion planning with T3Former.
Given the predicted scene changes, $\Delta\hat{s}^{t+k}$, provided by T3Former, we can generate a comprehensive motion plan by anticipating potential dangers based on both the scene changes and the previous and next scenes.
To achieve this, we first map the previous frame triplane $\hat{s}^{t+k-1}$, the triplane change $\Delta\hat{s}^{t+k}$, and the next frame triplane $\hat{s}^{t+k}$ through the ResNet-18 networks [11] $\theta_{p}$, $\theta_{\Delta}$, and $\theta_{f}$, respectively, into a shared latent space, denoted as $z^{t+k-1}$, $\Delta z^{t+k}$, and $z^{t+k}$. Then, $\Delta z^{t+k}$ passes through two independent fully connected (FC) layers and Sigmoid layers, collectively referred to as $\zeta_{p}$ and $\zeta_{f}$, to generate the query vectors $f^{t+k-1}$ and $f^{t+k}$, which are then multiplied with $z^{t+k-1}$ and $z^{t+k}$ to obtain the motion-related features of the scene change relative to the last triplane and the next triplane. Finally, the motion-related features and the scene change $\Delta z^{t+k}$ are element-wise averaged and processed through a FC layer, and then added to the future position encoding, resulting in the query variable for the next timestep motion plan, $Q^{t+k}_{a}\in\mathbb{R}^{d_{\rm act}}$.
Next, we project the actions of the historical $\tau_{p}$ frames into the same dimension as $Q^{t+k}_{a}$, and employ a Transformer encoder to capture the motion dependencies within these frames. Then, $Q^{t+k}_{a}$ is used in the Transformer decoder via cross-attention to predict the next action $\hat{a}^{t+k+1}$. The autoregressive generation of motion plans is defined as:
$\hat{a}^{t+k+1}=\bm{\Phi}_{{\rm act}}(\hat{a}^{t+k-\tau_{p}:t+k},\hat{s}^{t+k-1},\Delta\hat{s}^{t+k},Q^{t+k}_{a}).$ (10)
Unlike OccWorld [58], we do not require an additional ego token to continuously track the agent’s motion trajectory. Instead, we can learn future motion directly from the scene changes and historical motion, as shown in Figure 5.
The optimization objective is defined as follows:
$J_{\rm act}=\mathbb{E}_{t,k}[\mathcal{L}_{\rm act}(\hat{a}^{t+k},a^{t+k})],$ (11)
where $\mathcal{L}_{\rm act}$ measures the L2 discrepancy between the predicted and GT trajectories.
4 Experiments
4.1 Experimental settings
Targets and evaluation metrics. OWMs focus on jointly modeling occupancy forecasting and motion planning. Following the conventions in [58, 51, 46], we use the past four frames (2 seconds) to predict the outcomes of the next six frames (3 seconds). We conduct two sets of experiments:
$\bullet$ To validate the accuracy of 4D occupancy forecasting, we use intersection over union (IoU) and mean IoU (mIoU) across all semantic classes to measure the predicted future occupancy under the Occ3D dataset [39].
$\bullet$ To assess the precision and safety of motion planning, we calculate the L2 distance (in meters) between the planned and GT trajectories, and the collision rate with traffic participants’ bounding boxes using the nuScenes dataset [3].
Implementation details. The dataset consists of 1,000 scenes, of which 850 are used for training and 100 for testing. Each scene contains up to 40 timesteps, with a sampling frequency of 2Hz. The dimensions of the occupancy data $S^{t}$ at each timestep are $16\times 200\times 200$. The pre-trained Triplane shape is $16\times 100\times 100$ with 8 channels. In $\bm{\Phi}_{\rm fut}$, predictions for each plane incorporate features at $V=5$ different scales, while the token dimension in $\bm{\Phi}_{\rm act}$ is $d_{\rm act}=50$. The objectives, $J_{\rm enc,dec}$, $J_{\rm fut}$, and $J_{\rm mot}$, are optimized using AdamW with a weight regularization factor of 0.01, an initial learning rate of 0.001, and cosine decay with a minimum learning rate of $10^{-6}$. We first pre-train $\bm{\Phi}_{\rm enc}$ and $\Phi_{\rm dec}$ with a batch size of 10, using random flip augmentation to obtain the triplane representations. Then, we train $\Phi_{\rm fut}$ and $\bm{\Phi}_{\rm act}$ with a batch size of 1, while fine-tuning $\bm{\Phi}_{\rm dec}$. All training and testing are performed on 4 RTX 4090 GPUs. More details are provided in the supplementary material.
Table 1: Testing performance comparison with SOTA methods on the 4D occupancy forecasting task. Best values in each metric are bolded. 0s refers to reconstruction accuracy, while 1s, 2s, and 3s denote future prediction accuracy. Avg. is the average of 1s, 2s, and 3s.
Models Input mIoU (%) $\uparrow$ IoU (%) $\uparrow$
0s 1s 2s 3s Avg. 0s 1s 2s 3s Avg.
Copy&Paste 3D-Occ 66.38 14.91 10.54 8.52 11.33 62.29 24.47 19.77 17.31 20.52
OccWorld-O [58] 3D-Occ 66.38 25.78 15.14 10.51 17.14 62.29 34.63 25.07 20.18 26.63
OccLLaMA-O [46] 3D-Occ 75.20 25.05 19.49 15.26 19.93 63.76 34.56 28.53 24.41 29.17
RenderWorld [53] 3D-Occ - 28.69 18.89 14.83 20.80 - 37.74 28.41 24.08 30.08
OccLLM-O [51] 3D-Occ - 24.02 21.65 17.29 20.99 - 36.65 32.14 28.77 32.52
DOME-O [8] 3D-Occ 83.08 35.11 25.89 20.29 27.10 77.25 43.99 35.36 29.74 36.36
T3Former-O 3D-Occ 85.50 46.32 33.23 28.73 36.09 92.07 77.00 75.89 76.32 76.40
OccWorld-F [58, 27] Camera 20.09 8.03 6.91 3.54 6.16 35.61 23.62 18.13 15.22 18.99
OccLLaMA-F [46, 27] Camera 37.38 10.34 8.66 6.98 8.66 38.92 25.81 23.19 19.97 22.99
RenderWorld [53] Camera - 2.83 2.55 2.37 2.58 - 14.61 13.61 12.98 13.73
OccLLM-F [51, 27] Camera - 11.28 10.21 9.13 10.21 - 27.11 24.07 20.19 23.79
DOME-F [8] Camera 75.00 24.12 17.41 13.24 18.25 74.31 35.18 27.90 23.44 28.84
T3Former-F [27] Camera 43.52 24.87 18.30 15.63 19.60 54.31 38.98 37.45 31.89 36.11
4.2 Comparisons with the state-of-the-art
4.2.1 4D occupancy forecasting
Table 1 presents the performance results of various methods, considering two modes of operation: (i) using 3D occupancy GTs as historical input, marked with “-O”; (ii) using predicted 3D occupancy data from FB-Occ [27] as historical input, marked with “-F”. “Copy&Paste” refers to directly using the GT occupancy data from the current timestep as predictions for future outcomes.
It is clear that SOTA methods’ $\bm{\Phi}_{\rm enc}$ fail to achieve high-fidelity compression, resulting in low reconstruction accuracy at time 0s. In contrast, both T3Former-O and T3Former-F utilize triplane representations that preserve both geometric and semantic information, leading to the highest mIoU and IoU reconstruction accuracy. This significantly aids in $\bm{\Phi}_{\rm fut}$ for predicting the next 3 seconds, effectively reducing error accumulation. Specifically, the “Copy&Paste” method, as a simple baseline, demonstrates substantial error accumulation. In contrast, T3Former-O shows a significant reduction in error, as indicated by mIoU and IoU averages. Notably, IoU generally surpasses mIoU, as predicting occupancy (whether occupied or unoccupied) is relatively straightforward, while semantic tracking without instance-level supervision is more challenging and prone to drift over time.
4.2.2 Motion planning
We compare T3Former extensively with SOTA methods for autonomous driving, including LiDAR-based (IL [34], NMP [54], FF [13], EO [22]), camera-based (ST-P3 [14], UniAD [16], VAD [19], OccNet [40]), and occupancy-based methods (OccNet [40], OccWorld [58], RenderWorld [53], OccLLaMA [46]). The results are shown in Table 2.
Table 2: Testing performance of motion planning compared with SOTA method. Best and second-best values in each metric are bolded and underlined, respectively. Auxiliary supervision refers to additional supervision signals beyond the GT trajectories.
Models Input Auxiliary supervision L2 (m) $\downarrow$ Collision rate (%) $\downarrow$
1s 2s 3s Avg. 1s 2s 3s Avg.
IL [34] LiDAR None 0.44 1.15 2.47 1.35 0.08 0.27 1.95 0.77
NMP [54] LiDAR Box+Motion 0.53 1.25 2.67 1.48 0.04 0.12 0.87 0.34
FF [13] LiDAR Freespace 0.55 1.20 2.54 1.43 0.06 0.17 1.07 0.43
EO [22] LiDAR Freespace 0.67 1.36 2.78 1.60 0.04 0.09 0.88 0.33
ST-P3 [14] Camera Map+Box+Depth 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71
UniAD [16] Camera Map+Box+Motion+Track+Occ 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31
VAD [19] Camera Map+Box+Motion 0.54 1.15 1.98 1.22 0.04 0.39 1.17 0.53
OccNet [40] Camera Map+Box+3D-Occ 1.29 2.13 2.99 2.14 0.21 0.59 1.37 0.72
OccNet [40] 3D-Occ Map+Box 1.29 2.31 2.98 2.25 0.20 0.56 1.30 0.69
OccWorld-O [58] 3D-Occ None 0.43 1.08 1.99 1.17 0.07 0.38 1.35 0.60
RenderWorld [53] 3D-Occ None 0.35 0.91 1.84 1.03 0.05 0.40 1.39 0.61
OccLLaMA-O [46] 3D-Occ None 0.37 1.02 2.03 1.14 0.04 0.24 1.20 0.49
T3Former-O 3D-Occ None 0.32 0.91 1.76 1.00 0.08 0.32 0.51 0.30
Specifically, LiDAR- and camera-based methods explore various combinations of auxiliary supervisions to enhance the quality of planned motion trajectories. Despite their success, the auxiliary information they rely on is labor-intensive to obtain. In contrast, occupancy-based methods generate motion proposals using only occupancy representations. T3Former achieves the lowest L2 error over time, the lowest collision rate at the final time step, and the lowest average collision rate across all future motions. Compared to SOTA occupancy-based methods, T3Former shows excellent short-term and long-term planning (0s and 3s). This is primarily because we can establish precise, task-relevant correlations between scene changes and motion trajectories in a deep latent space, effectively filtering out noise from scene variations.
Figure 6: Visualization of 4D occupancy forecasting and motion planning for the next 3 seconds on the Occ3D test dataset (zoom in for a clearer view).
4.3 Visualizations
Comparisons with SOTA.
Figure 6 illustrates the evolution of the scene over the next 3 seconds, along with the corresponding motion predictions for two different scenarios. (i) In the first case, it is clear that OccWorld-O incorrectly labels the truck as a trailer initially, resulting in a continuous accumulation of errors. In contrast, T3Former effectively learns the motion dynamics of all objects in the scene, consistently tracking each vehicle over a long horizon, while providing clearer object boundaries. (ii) In the second
Figure 7: Driving view generations conditioned on different weather prompts and different triplane constraints.
Figure 8: Predicted locations of dynamic objects for the three last frames.
case, as the prediction horizon extends, OccWorld-O’s road boundary gradually fades, losing its clarity, whereas our method maintains the boundary shape of the road, including the trees. Additionally, OccWorld-O misidentifies the motorcycle as a truck at both 2.5s and 3s, while T3Former-O ensures that the semantics of small objects remain accurate over time. (iii) Under both scenarios, T3Former-O enables precise long-term future predictions, allowing our motion planning to rely solely on the evolving scene dynamics. This leads to more accurate trajectories and better avoidance of potential future hazards.
Driving view generation. Recent diffusion-based autonomous driving models [44, 7, 45] typically use historical driving views as conditions to generate future views. In contrast, our approach uses only triplane data as a lightweight condition. We fine-tune the diffusion model with ControlNet [56]. Figure 8 showcases synthetic images generated with different weather prompts and different triplane conditions. Triplane provides accurate 3D structural information, thereby enhancing consistency between pixel positions and semantics in the resulting synthetic 2D images.
Comparison of motion prediction. Figure 8 compares occupancy predictions for the three furthest frames in the prediction window, focusing on dynamic objects of varying sizes (cars and pedestrians) from a BEV perspective. Two reference lines are displayed to reflect
Figure 9: Limitations in sparse boundary regions.
Table 3: Ablation study of T3Former.
Idx. Models Avg. mIoU $\uparrow$ Avg. L2 $\downarrow$ FPS $\uparrow$
M0 T3Former-O 36.09 1.00 26
M1 w/o pretraining 33.96 1.17 26
M2 w/o triplane 28.97 1.10 21
M3 w/o triplane changes 30.51 1.04 27
M4 w/o multi-scale mot. 31.92 1.12 36
M5 w/o autoregression 34.65 1.06 34
Table 4: Effect of different hyperparameters.
Idx. Setting Avg. mIoU $\uparrow$ Avg. L2 $\downarrow$ FPS $\uparrow$
H1.1 8, (8, 50, 50) 26.78 1.33 37
H1.2 8, (16, 50, 50) 28.91 1.23 33
H1.3 8, (8, 100, 100) 31.42 1.17 30
H1.4 8, (16, 100, 100) 36.09 1.00 26
H1.5 16, (16, 100, 100) 36.20 1.09 23
H1.6 32, (16, 100, 100) 35.71 1.11 18
H2.1 $V=3$, (4, 16) 35.91 1.08 28
H2.2 $V=5$, (4, 16) 36.09 1.00 26
H2.3 $V=5$, (4, 32) 36.14 1.04 21
H2.4 $V=5$, (8, 16) 36.18 1.02 23
H2.5 $V=7$, (4, 16) 35.65 0.98 20
the errors between the predictions and GT. The results of our method match the GT significantly better than OccWorld-O, which exhibits substantial drift. For instance, the predictions for the pedestrians at the top and the vehicle on the left clearly demonstrate this improvement
Limitation. As shown in Figure 9, T3Former produces sparse predictions in boundary regions. This is primarily due to occlusions in the original occupancy data, coupled with the use of triplanes that employ the orthogonal decomposition strategy [25, 17] to encode fine-grained geometric details. While this approach is effective for detailed representation, it complicates the task of correlating surrounding context and filling in the sparse areas during forecasting. In contrast, OccWorld tends to smooth and homogenize these sparse regions, a result of the inherent blurring issue in VAE-based models [5, 12]. In future work, we aim to improve content generation in these sparsely observed regions.
4.4 Ablation study
Table 3 validates the effectiveness of the key designs in T3Former. M0 represents our complete version. For fairness, all ablation models are derived by replacing specific modules in M0, as outlined below: M1 directly learns triplanes and downstream tasks (Eqs. (9) and (11)) in an end-to-end manner, without pre-training, leading to reduced encoder efficiency and performance degradation. M2 replaces T3Former’ input with the VQ-VAE token from OccWorld [58], causing some geometric information loss. M3 predicts the entire future triplane instead of $\Delta$ changes, showing that change prediction is more efficient. M4 uses a single-scale Transformer to model temporal dynamics, failing to capture multi-scale object motion and causing error accumulation. M5 predicts all future outcomes in parallel, resulting in poor performance due to the inability to correct autoregressive errors.
Table 4 analyzes the effect of different hyperparameters. denotes our trade-off version. The H1 series illustrates the impact of different latent shapes of triplanes. The setting refers to the number of channels, as well as the (height, width, and length). It is clear that increasing the spatial shape significantly enhances future prediction performance, while larger channel numbers provide little benefit and may even reduce performance. This is due to the introduction of feature redundancy, which complicates self-attention learning during forecasting. In the H2 series, the setting defines the number of scales, Transformer layers, and self-attention heads in T3Former. Our results show that even a lightweight Transformer effectively captures key patterns, while increasing model size brings little benefit but adds computational overhead. Consequently, our design uses only 55M parameters – 23% fewer than OccWorld’s 72M – yet achieves superior efficiency. As illustrated in Figure 2, T3Former runs at 26 FPS on an RTX 4090 GPU, outperforming OccWorld’s 18 FPS.
5 Conclusion
This paper presents T3Former, a new 4D occupancy world model with three key designs: (i) pre-trained triplane representations for OWM, (ii) multi-scale Transformers for capturing motion dynamics, and (iii) triplane change prediction for occupancy forecasting and motion planning. T3Former achieves SOTA performance in 4D occupancy forecasting and motion planning with real-time capability. Our future work will dive deeper into driving view generation (see Figure 8), focusing on the synergy between prompt constraints, image texture, and temporal consistency in occupancy data.